# A Common Origin of Asymmetric Self-interacting Dark Matter and Dirac Leptogenesis

### v1

@@ Abstract
Assuming dark matter to be asymmetric as well as self-interacting and neutrinos to be Dirac fermions, we propose a framework to address the observed baryon imbalance of the universe. We add three right-handed neutrinos $\nu_{R_{i}},\,{i=1,2,3}$, one singlet fermion $\chi$, a doublet fermion $\psi$, and heavy scalar doublets $\eta_{i},\,{i=1,2}$ to the Standard Model. Both $\chi$ and $\psi$ are fermions with non-zero charge under an extended $U(1)_{B-L}\times U(1)_{D}$ symmetry. Additionally, a $\mathcal{Z}_{2}$ symmetry is imposed, where the singlets $\chi$, $\nu_{R}$, and $\eta$ are negative and the doublet $\psi$ is positive. Given the assumption that neutrinos are Dirac particles, $B-L$ turns into an exact symmetry of the universe. The CP-violating out-of-equilibrium decay of heavy scalar $\eta$ generates an equal and opposite $B-L$ asymmetry among the left-handed ($\nu_{L}$) and right-handed ($\nu_{R}$) neutrinos. The $\nu_{L}-\nu_{R}$ equilibration process does not take place until below the Electroweak phase transition scale because of tiny Yukawa couplings. During this time, Sphaleron processes, which are active at temperatures higher than 100 GeV, transform a portion of the $B-L$ asymmetry stored in left-handed neutrinos into baryon asymmetry. MeV scale gauge boson $Z^{\prime}$ of $U(1)_{D}$ sector mediates both annihilation of symmetric dark matter component and self-interaction among dark matter particles. Moreover, $Z^{\prime}$ mixes with the Standard Model Z-boson and provides a portal for dark matter direct detection.
@@ I Introduction
where, $n_{B}$, $n_{\overline{B}}$ and $n_{\gamma}$ are the number densities of baryon, antibaryon and photon respectively. Based on Sakharov conditions Sakharov:1967dj ; sakharov.67 , the observed baryon asymmetry has been addressed in frameworks of leptogenesis fukugita.86 ; baryo_lepto_group_1 ; baryo_lepto_group_2 ; baryo_lepto_group_3 ; baryo_lepto_group_4 ; baryo_lepto_group_5 ; baryo_lepto_group_6 ; baryo_lepto_group_7 ; baryo_lepto_group_8 ; baryo_lepto_group_9 ; Abada:2006ea ; Pilaftsis:2003gt ; Pilaftsis:2005rv ; Kaplan:2009ag ; Feng:2012jn ; Kolb:1990vq ; Ibe:2011hq ; Feng:2013wn ; Narendra:2018vfw . WMAP and Planck data imply that $\Omega_{\rm DM}\approx 5\Omega_{\rm B}$. It triggers to assume DM to have a similar origin as that of baryons and address both phenomenologies in a common framework old_DM_Baryon_asy_1 ; old_DM_Baryon_asy_2 ; old_DM_Baryon_asy_3 ; old_DM_Baryon_asy_4 ; old_DM_Baryon_asy_5 ; old_DM_Baryon_asy_6 ; old_DM_Baryon_asy_7 ; old_DM_Baryon_asy_8 ; Asydm_models1_1 ; Asydm_models1_2 ; Asydm_models1_3 ; Asydm_models1_4 ; Asydm_models1_5 ; Asydm_models1_6 ; Asydm_models1_7 ; Asydm_models1_8 ; Asydm_models1_9 ; Asydm_models1_10 ; Asydm_models2_1 ; Asydm_models2_2 ; Asydm_models2_3 ; Asydm_models2_4 ; Asydm_models2_5 ; Asydm_models2_6 ; Asydm_models2_7 ; Asydm_models2_8 ; Asydm_models2_9 ; Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl ; Asydm_review_1 ; Asydm_review_2 .
@@
From cosmological standpoint, DM is assumed to be a cold, collisionless fluid by the ${\rm\Lambda CDM}$ model. ${\rm\Lambda CDM}$ model has proved incredibly successful in explaining the large-scale structures of the Universe. Nevertheless, a number of discrepancies have surfaced at smaller scales, such as the core-cusp problem, the missing satellite problem, and the too-big-to-fail problem etc Tulin:2017ara ; Bullock:2017xww . These discrepancies are more noticeable at the dwarf galaxy scale and progressively diminish to the cluster scale, which corresponds with $\Lambda{\rm CDM}$ predictions at large scales. In order to mitigate these anomalies, self-interacting dark matter (SIDM) Carlson:1992fn ; deLaix:1995vi ; Spergel:1999mh was suggested as a substitute for traditional cold dark matter. A velocity-dependent self-scattering of order $\sigma/m\sim 1\;{\rm cm}^{2}/{\rm g}\approx 2\times 10^{-24}\;{\rm cm}^{2}/{\rm GeV}$ has been demonstrated to alleviate these anomalies Buckley:2009in ; Feng:2009hw ; Feng:2009mn ; Loeb:2010gj ; Zavala:2012us ; Vogelsberger:2012ku ; Bringmann:2016din ; Kaplinghat:2015aga ; Kamada:2020buc ; Aarssen:2012fx ; Tulin:2013teo . DM models with a light mediator is well suited to realise such velocity dependence Kouvaris:2014uoa ; Bernal:2015ova ; Kainulainen:2015sva ; Hambye:2019tjt ; Cirelli:2016rnw ; Kahlhoefer:2017umn ; Ghosh:2021wrk . However a common issue with SIDM is its under-abundant relic due to efficient annihilation into the light mediator Dutta:2021wbn ; Borah:2021yek ; Borah:2021pet ; Borah:2021rbx ; Dutta:2024ydz . However, if SIDM is assumed to have similar origin as that of baryons, correct relic density can be achieved Dutta:2022knf ; Borah:2024wos .
@@
Considering that neutrinos are Dirac particles (i.e., $B-L$ is an exact symmetry), Dirac Leptogenesis can account for the observed baryon asymmetry of the Universe Dirac_leptogenesis_1 ; Dirac_leptogenesis_2 ; Dirac_leptogenesis_3 ; Dirac_leptogenesis_4 ; Dirac_leptogenesis_5 ; Dirac_leptogenesis_6 ; Dirac_leptogenesis_7 . The fundamental mechanism of Dirac Leptogenesis is that the equilibration time between left and right-handed neutrinos mediated via Yukawa interaction with SM Higgs (i.e. $y\overline{\nu_{R}}H\nu_{L}$) is much less than the $(B+L)$ violating sphaleron transitions above electroweak phase transition. Therefore, demanding $B-L=B-(L_{\rm SM}+L_{\nu_{R}})=0$Cerdeno:2006ha , a net $B-L_{\rm SM}$ can be generated in terms of $L_{\nu_{R}}$. $\nu_{R}$ being $SU(2)_{L}$ singlet, $L_{\nu_{R}}$ remains unaffected by sphalerons. The non-zero $B-L_{\rm SM}$ is, however, converted to a net $B$ asymmetry via $B+L$ violating sphaleron transitions Dick ; Buchmuller:2000as ; Cerdeno:2006ha ; Gu:2006dc ; Gu:2007mc ; Murayama:2002je ; Thomas:2005rs ; Borah:2016zbd ; Perez:2021udy .
To integrate all these avenues, in this paper, we construct a framework for explaining DM and baryon asymmetry considering Dirac neutrinos. The asymmetry in both sectors is generated by the CP-violating out-of-equilibrium decay of a heavy scalar doublet $\eta$ into $\ell\nu_{R}$ and $\psi\chi$, where $\ell$ is SM lepton, $\nu_{R}$ is right handed neutrino (RHN), and $\psi$, $\chi$ are respectively a vector-like fermion doublet and a vector-like fermion singlet that comprises the dark sector. The lepton asymmetry in the channel $\eta\to\ell\nu_{R}$ is then converted to a net baryon asymmetry via $B+L$ violating sphaleron transitions. Being a vector-like Dirac fermion, asymmetry in the channel $\eta\to\psi\chi$ remains intact Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl . The dark sector is gauged under a $U(1)_{D}$ symmetry. The corresponding gauge boson $Z^{\prime}$ mediates DM self interaction. The lightest dark sector particle $\chi$ which is odd under $\mathcal{Z}_{2}$ acts as the SIDM candidate. The symmetric DM component annihilates away into a pair of $Z^{\prime}$, yielding a vanishingly small symmetric relic. We also impose $U(1)_{B-L}\times\mathcal{Z}_{2}$ symmetry which ensures that the neutrinos are Dirac. Meanwhile, $\mathcal{Z}_{2}$ symmetry is softly broken by a term $\mu^{2}\eta^{\dagger}H$, generating Dirac mass for neutrinos. A schematic of the framework is shown in Fig. 1.
@@
The paper is organized as follows. In sec. II, we discuss the proposed model, while in sec. III, we explain the Dirac mass of light neutrinos. Section IV is devoted to explain baryogenesis via leptogenesis from the decay of heavy particles $\eta$, while section V describes the mechanism for DM relic abundance. The phenomenology of SIDM is discussed in section VI, followed by constraints from direct search in sec. VII. We finally conclude in sec. VIII.
@@ II The Model
We extend the SM both by new symmetries and particles. The new particles include: three RHNs $\nu_{{R}_{\alpha}}$, $\alpha=1,2,3$, a heavy scalar doublet $\eta$, and a pair of vector-like leptons, one singlet ($\chi$) and the other doublet ($\psi$) under $SU(2)$. A discrete symmetry $\mathcal{Z}_{2}$ is introduced under which $\eta,\nu_{R}$ and $\chi$ are odd, while $\psi$ and all other SM particles are even. The fermions $\psi$ and $\chi$ also carry quantum numbers under $U(1)_{D}$ symmetry. $\chi$ is the lightest of dark sector particles, hence the DM candidate. $U(1)_{D}$ symmetry is spontaneously broken and MeV scale $U(1)_{D}$ gauge boson $Z^{\prime}$ mediates DM self-interaction. The overall $B-L$ global symmetry remains intact because neutrinos are considered to be Dirac fermions. The particle content beyond those in the SM and their charge assignments are shown in Table 1.
@@
Heavy scalar $\eta$ undergoes CP-violating out-of-equilibrium decay. This decay creates asymmetries in both lepton and dark sectors simultaneously Arina:2011cu . The process $\eta\to\ell\nu_{R}$, creates equal and opposite lepton asymmetry in left and right-handed sectors. $B+L$ violating sphaleron transitions converts the lepton asymmetry in the left-handed sector to a net baryon asymmetry. Asymmetry in the right-handed sector remains unaffected until the temperature falls much below the electroweak phase transition. DM $\chi$ attains thermal equilibrium in early universe via processes like $\psi\psi\to\chi\chi$, mediated by $Z^{\prime}$. However, $\chi$ annihilates efficiently into pairs of $Z^{\prime}$ via t-channel and a negligible symmetric DM component is left. The correct DM relic density is obtained from the CP-violating out-of-equilibrium decay of heavy scalar $\eta$. The relevant Lagrangian is given by:
$\begin{split}\mathcal{L}&\supset\bar{\psi}i\gamma^{\mu}D_{\mu}\psi+\bar{\chi}i\gamma^{\mu}D^{\prime}_{\mu}\chi+M_{\psi}\bar{\psi}\psi+M_{\chi}\bar{\chi}\chi\\ &+\left[f_{kl}\overline{\ell_{k}}\tilde{\eta}{\nu_{R}}_{l}+\lambda\overline{\psi}\tilde{\eta}\chi+{\rm h.c.}\right]-V(H,\eta)\,,\end{split}$ (2)
@@
The gauge boson $Z^{\prime}$ of the spontaneously broken $U(1)_{D}$ symmetry facilitates DM self-interaction. Hence, in this model, the observed DM relic is dominantly asymmetric. Kinetic mixing between $Z^{\prime}$ and SM $Z$ enables direct detection of DM in terrestrial laboratories. We elaborate these phenomenologies in respective sections.

### v2

@@ Abstract
Assuming dark matter to be asymmetric as well as self-interacting and neutrinos to be Dirac fermions, we propose a framework to address the observed baryon imbalance of the universe. We add three right-handed neutrinos $\nu_{R_{i}},\,{i=1,2,3}$, one singlet fermion $\chi$, a doublet fermion $\psi$, and heavy scalar doublets $\eta_{i},\,{i=1,2}$ to the Standard Model. A global BLB-L is imposed to protect the Dirac nature of neutrinos. Both $\chi$ and $\psi$ are fermions with non-zero charge under an extended $U(1)_{B-L}\times U(1)_{D}$ symmetry. Additionally, a $\mathcal{Z}_{2}$ symmetry is imposed, where the singlets $\chi$, $\nu_{R}$, and $\eta$ are negative and the doublet $\psi$ is positive. The CP-violating out-of-equilibrium decay of heavy scalar $\eta$ generates an equal and opposite BLB-L asymmetry among the left-handed ($\nu_{L}$) and right-handed ($\nu_{R}$) neutrinos. The $\nu_{L}-\nu_{R}$ equilibration process does not take place until below the Electroweak phase transition scale because of tiny Yukawa couplings. During this time, Sphaleron processes, which are active at temperatures higher than 100 GeV, transform a portion of the BLB-L asymmetry stored in left-handed neutrinos into baryon asymmetry. MeV scale gauge boson $Z^{\prime}$ of $U(1)_{D}$ sector mediates both annihilation of symmetric dark matter component and self-interaction among dark matter particles. Moreover, $Z^{\prime}$ mixes with the Standard Model Z-boson and provides a portal for dark matter direct detection.
@@ I Introduction
where, $n_{B}$, $n_{\overline{B}}$ and $n_{\gamma}$ are the number densities of baryon, antibaryon and photon respectively. Based on Sakharov conditions Sakharov:1967dj ; sakharov.67 , the observed baryon asymmetry has been addressed in frameworks of leptogenesis fukugita.86 ; baryo_lepto_group_1 ; baryo_lepto_group_2 ; baryo_lepto_group_3 ; baryo_lepto_group_4 ; baryo_lepto_group_5 ; baryo_lepto_group_6 ; baryo_lepto_group_7 ; baryo_lepto_group_8 ; baryo_lepto_group_9 ; Abada:2006ea ; Pilaftsis:2003gt ; Pilaftsis:2005rv ; Kaplan:2009ag ; Feng:2012jn ; Kolb:1990vq ; Ibe:2011hq ; Feng:2013wn ; Narendra:2018vfw . WMAP and Planck data imply that $\Omega_{\rm DM}\approx 5\Omega_{\rm B}$. It triggers to assume DM to have a similar origin as that of baryons and address both phenomenologies in a common framework old_DM_Baryon_asy_1 ; old_DM_Baryon_asy_2 ; old_DM_Baryon_asy_3 ; old_DM_Baryon_asy_4 ; old_DM_Baryon_asy_5 ; old_DM_Baryon_asy_6 ; old_DM_Baryon_asy_7 ; old_DM_Baryon_asy_8 ; Asydm_models1_1 ; Asydm_models1_2 ; Asydm_models1_3 ; Asydm_models1_4 ; Asydm_models1_5 ; Asydm_models1_6 ; Asydm_models1_7 ; Asydm_models1_8 ; Asydm_models1_9 ; Asydm_models1_10 ; Asydm_models2_1 ; Asydm_models2_2 ; Asydm_models2_3 ; Asydm_models2_4 ; Asydm_models2_5 ; Asydm_models2_6 ; Asydm_models2_7 ; Asydm_models2_8 ; Asydm_models2_9 ; Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl ; Asydm_review_1 ; Asydm_review_2 ; Kanemura:2025ixp .
@@
From cosmological standpoint, DM is assumed to be a cold, collisionless fluid by the $\rm\Lambda CDM$} model. $\rm\Lambda CDM$} model has proved incredibly successful in explaining the large-scale structures of the Universe. Nevertheless, a number of discrepancies have surfaced at smaller scales, such as the core-cusp problem, the missing satellite problem, and the too-big-to-fail problem etc Tulin:2017ara ; Bullock:2017xww . These discrepancies are more noticeable at the dwarf galaxy scale and progressively diminish to the cluster scale, which corresponds with $\Lambda{\rm CDM}$ predictions at large scales. In order to mitigate these anomalies, self-interacting dark matter (SIDM) Carlson:1992fn ; deLaix:1995vi ; Spergel:1999mh was suggested as a substitute for traditional cold dark matter. A velocity-dependent self-scattering of order $\sigma/m\sim 1\;{\rm cm}^{2}/{\rm g}\approx 2\times 10^{-24}\;{\rm cm}^{2}/{\rm GeV}$ has been demonstrated to alleviate these anomalies Buckley:2009in ; Feng:2009hw ; Feng:2009mn ; Loeb:2010gj ; Zavala:2012us ; Vogelsberger:2012ku ; Bringmann:2016din ; Kaplinghat:2015aga ; Kamada:2020buc ; Aarssen:2012fx ; Tulin:2013teo . DM models with a light mediator is well suited to realise such velocity dependence Kouvaris:2014uoa ; Bernal:2015ova ; Kainulainen:2015sva ; Hambye:2019tjt ; Cirelli:2016rnw ; Kahlhoefer:2017umn ; Ghosh:2021wrk ; Yang:2025dgl . However a common issue with SIDM is its under-abundant relic due to efficient annihilation into the light mediator Dutta:2021wbn ; Borah:2021yek ; Borah:2021pet ; Borah:2021rbx ; Dutta:2024ydz . However, if SIDM is assumed to have similar origin as that of baryons, correct relic density can be achieved Dutta:2022knf ; Borah:2024wos .
@@
Considering that neutrinos are Dirac particles (which requires BLB-L to be an exact symmetry), Dirac Leptogenesis can account for the observed baryon asymmetry of the Universe Dirac_leptogenesis_1 ; Dirac_leptogenesis_2 ; Dirac_leptogenesis_3 ; Dirac_leptogenesis_4 ; Dirac_leptogenesis_5 ; Dirac_leptogenesis_6 ; Dirac_leptogenesis_7 . The fundamental mechanism of Dirac Leptogenesis is that the equilibration time between left and right-handed neutrinos mediated via Yukawa interaction with SM Higgs (i.e. $y\overline{\nu_{R}}H\nu_{L}$) is much less than the $(B+L)$ violating sphaleron transitions above electroweak phase transition. Therefore, demanding $B-L=B-(L_{\rm SM}+L_{\nu_{R}})=0$Cerdeno:2006ha , a net $B-L_{\rm SM}$ can be generated in terms of $L_{\nu_{R}}$. $\nu_{R}$ being $SU(2)_{L}$ singlet, $L_{\nu_{R}}$ remains unaffected by sphalerons. The non-zero $B-L_{\rm SM}$ is, however, converted to a net $B$ asymmetry via $B+L$ violating sphaleron transitions Dick ; Buchmuller:2000as ; Cerdeno:2006ha ; Gu:2006dc ; Gu:2007mc ; Murayama:2002je ; Thomas:2005rs ; Borah:2016zbd ; Perez:2021udy .
To integrate all these avenues, in this paper, we construct a framework for explaining DM and baryon asymmetry considering Dirac neutrinos. The asymmetry in both sectors is generated by the CP-violating out-of-equilibrium decay of a heavy scalar doublet $\eta$ into $\ell\nu_{R}$ and $\psi\chi$, where $\ell$ is SM lepton, $\nu_{R}$ is right handed neutrino (RHN), and $\psi$, $\chi$ are respectively a vector-like fermion doublet and a vector-like fermion singlet that comprises the dark sector. The lepton asymmetry in the channel $\eta\to\ell\nu_{R}$ is then converted to a net baryon asymmetry via $B+L$ violating sphaleron transitions. Being a vector-like Dirac fermion, asymmetry in the channel $\eta\to\psi\chi$ remains intact Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl . The dark sector is gauged under a $U(1)_{D}$ symmetry. The corresponding gauge boson $Z^{\prime}$ mediates DM self interaction. The lightest dark sector particle $\chi$ which is odd under $\mathcal{Z}_{2}$ acts as the SIDM candidate. The symmetric DM component annihilates away into a pair of $Z^{\prime}$, yielding a vanishingly small symmetric relic. We also impose $U(1)_{B-L}\times\mathcal{Z}_{2}$ symmetry which ensures that the neutrinos are Dirac. $\mathcal{Z}_{2}$ symmetry is softly broken by a term $\mu^{2}\eta^{\dagger}H$, generating Dirac mass for neutrinos. A schematic of the framework is shown in Fig. 1.
@@
The paper is organized as follows. In sec. II, we discuss the proposed model, while in sec. III, we explain the Dirac mass of light neutrinos. Section IV is devoted to explain baryogenesis via leptogenesis from the decay of heavy particles $\eta$, while section V describes the mechanism for DM relic abundance. The phenomenology of SIDM is discussed in section VI, followed by constraints from direct search in sec. VII. Then we discuss possible collider phenomenology in sec. VIII. We finally conclude in sec. IX.
@@ II The Model
We extend the SM both by new symmetries and particles. The new particles include: three RHNs $\nu_{{R}_{\alpha}}$, $\alpha=1,2,3$, a heavy scalar doublet $\eta$, and a pair of vector-like leptons, one singlet ($\chi$) and the other doublet ($\psi$) under $SU(2)$. The particle content beyond those in the SM and their charge assignments under the extended symmetries are shown in Table 1.
@@
A discrete symmetry $\mathcal{Z}_{2}$ is introduced under which $\eta,\nu_{R}$ and $\chi$ are odd, while $\psi$ and all other SM particles are even. The fermions $\psi$ and $\chi$ also carry quantum numbers under $U(1)_{D}$ symmetry. $\chi$ is the lightest of dark sector particles, hence the DM candidate. $U(1)_{D}$ symmetry is spontaneously broken and MeV scale $U(1)_{D}$ gauge boson $Z^{\prime}$ mediates DM self-interaction. A global $U(1)_{B-L}$ symmetry is imposed so that only Dirac mass term is allowed for neutrinos, while Majorana mass terms are forbidden, ensuring lepton number conservation at the level of the renormalisable Lagrangian. Owing to charge assignments of the fields as in Table 1, the relevant Lagrangian for the phenomenologies is given by:
[…]
@@
Heavy scalar $\eta$ undergoes CP-violating out-of-equilibrium decay. This decay creates asymmetries in both lepton and dark sectors simultaneously Arina:2011cu . The process $\eta\to\ell\nu_{R}$, creates equal and opposite lepton asymmetry in left and right-handed sectors. $B+L$ violating sphaleron transitions converts the lepton asymmetry in the left-handed sector to a net baryon asymmetry. Asymmetry in the right-handed sector remains unaffected until the temperature falls much below the electroweak phase transition. DM $\chi$ attains thermal equilibrium in early universe via processes like $\psi\psi\to\chi\chi$, mediated by the $U(1)_{D}$ gauge boson $Z^{\prime}$. However, $\chi$ annihilates efficiently into pairs of $Z^{\prime}$ via t-channel and a negligible symmetric DM component is left. The correct DM relic density is obtained from the CP-violating out-of-equilibrium decay of heavy scalar $\eta$. Hence, in this model, the observed DM relic is dominantly asymmetric. Kinetic mixing between $Z^{\prime}$ and SM $Z$ enables direct detection of DM in terrestrial laboratories. $Z^{\prime}$ also facilitates DM self-interaction. The term $\mu^{2}\eta^{\dagger}H$ breaks $\mathcal{Z}_{2}$ symmetry softly generating Dirac mass for neutrinos. The global BLB-L symmetry ensures Dirac nature of the neutrinos forbidding Majorana mass.
However, it should be noted that, global symmetries are generically expected to be violated by quantum-gravitational effects Banks:2010zn . Consequently, a global $U(1)_{B-L}$, even if imposed at the renormalisable level, may be broken by Planck-suppressed operators that reintroduce Majorana mass terms. For this reason, the theoretically robust way to protect Dirac neutrino masses is to gauge the $U(1)_{B-L}$ symmetry. In a gauged theory, Majorana mass terms are forbidden by gauge invariance at all orders, and quantum-gravity effects cannot violate the symmetry since possible Planck-suppressed operators are tightly constrained by gauge charge assignments. With the charges listed in Table 1, our Lagrangian is already invariant under $U(1)_{B-L}$, and the field content (SM plus three $\nu_{R}$) ensures anomaly cancellation. Dark fermions $\psi$ and $\chi$ being vectorlike, do not contribute to gauge anomalies. Thus, the model admits a consistent gauged BLB-L embedding without altering interactions relevant to the Dirac leptogenesis mechanism. Gauging BLB-L simply introduces the corresponding gauge boson together with the usual kinetic term and covariant derivatives, plus an optional BLB-L–breaking scalar $\phi_{B-L}$ whose BLB-L charge can be chosen so as not to generate the unwanted operator $\phi_{B-L}\nu_{R}\nu_{R}$, thereby preserving the Dirac nature of neutrinos. If the BLB-L breaking scale is sufficiently high ($M_{Z_{B-L}}\gg\mathrm{TeV}$), the new gauge boson effectively decouples, leaving the low-energy phenomenology of our model unchanged. In this case, the primary role of gauged BLB-L is to protect the Dirac neutrino mass from quantum-gravity violations of global symmetries.